OpenStudy (anonymous):

All 28 members in Crestview High School’s Ski Club went on a one-day ski trip. Members can rent skis for $16.00 per day or snowboards for $19.00 per day. The club paid a total of $478 for rental equipment. How many members rented skis and how many rented snowboards?

OpenStudy (anonymous):

The way I would approach such a question is to put it into equations. My first equation would be 16x+19y=478 which represents $16 on ski's per person and $19 on snowboards must add up to $478 My second would be x+y=28 because the number of skis and snowboards rented must add up to 28 so each person has one item. Next you solve them simultaniously. Gonna use elimination 16x+19y=478 (1) x+y=28 (2) 16x+16y=448 (3 = equation 2 * 16) (1)-(3) => 3y=30 y = 10. So now we know that 10 people rented snow boards. Just sub that into eq (2) and get x = 18. So 10 people got snow boards and 18 ski's. To check, 10x19 = 190 16x18= 228 190+228=478 which works out perfectly!